OpenStudy (lordhelix8th):

A bag contains 5 red cubes and 7 white cubes. Celinda will select one cube, put it in her pocket, and then select a second cube. Which expression gives the probability that she will select two red cubes? A. B. C. D.

OpenStudy (lordhelix8th):

a= 5/12*5/12 b=5/12*4/12 c=5/12*5/11 d=5/12*4/11

OpenStudy (lordhelix8th):

@welshfella @MattAlva98 @samanthagreer @Australopithecus

OpenStudy (anonymous):

i think its b i could be wrong

OpenStudy (anonymous):

Ok so she has a total of 12 cubes: 5 red and 7 white cubes So when she firsts picks out a cube the probability that she will get a red cube is 5/12 She takes that red cube and places it in her pocket Now there are only 11 cubes left which 4 of them are red and 7 are white So the probability that she will choose a red again is 4/11

OpenStudy (anonymous):

so using the information Which expression best represents this scenario?

OpenStudy (lordhelix8th):

b @samjordon

OpenStudy (anonymous):

Well why would u say b? Read my explanation again

OpenStudy (anonymous):

but theres 12 cubes together

OpenStudy (lordhelix8th):

sorry d

OpenStudy (lordhelix8th):

i read it wrong

OpenStudy (lordhelix8th):

@samjordon its d

OpenStudy (anonymous):

Yea

OpenStudy (lordhelix8th):

Two baseball teams, the Jayhawks and the Falcons, each recorded the number of home runs hit in 1 season by every player on the team. Each team then calculated the mean number of home runs and the MAD (mean absolute deviation). Team Home runs Mean MAD Jayhawks 7.8 4.1 Falcons 8.2 3.8 Which statement best describes the overlap in the distribution of the two data sets? A. The overlap is high because both means are greater than either MAD. B. The overlap is high because the difference in the means is small compared to either MAD. C. The overlap is low because the difference in the MADs is small. D. The overlap is low because the sum of the means is large compared to either MAD.

OpenStudy (lordhelix8th):

@JackofallTradez

OpenStudy (lordhelix8th):

@SithsAndGiggles @shinebrightlikeadimon @andrewhaze

OpenStudy (lordhelix8th):

@Pocarii @PurplePanda312

OpenStudy (lordhelix8th):

@kidrah69

OpenStudy (lordhelix8th):

this one i think b